If you are given 237.74 g of the compound Lead (II) nitrate, keep answers to two decimal places.


How many moles would that contain?


How many particles would be present?


How many nitrate ions would be present?


1
Expert's answer
2021-03-11T08:47:25-0500

Mole = mass / molar mass =237.74/331g = 0.72 mole

No of particles = number of mole × avogadro constant

= 0.72 × 6.02×1023 particles = 4.33 × 1023 particles

There are 2 moles of nitrate ions in 1 mole of Pb(NO3)2

Therefore, = 2 × 4.33 ×1023 Ions = 8.66 ×1023 ions